OpenStudy (wiggler):

limit as x tends to infinity of sqrt(x^(1/x))

OpenStudy (wiggler):

\[\lim_{x \rightarrow \infty} \sqrt{x^\frac{ 1 }{ x }}\]Show working.

OpenStudy (anonymous):

Suppose \[L=\lim_{x\to\infty}\sqrt{x^{1/x}}\] Then \[\ln L=\ln\lim_{x\to\infty}\sqrt{x^{1/x}}~~\Rightarrow~~\ln L=\lim_{x\to\infty}\ln\sqrt{x^{1/x}}\] Since \(\sqrt{x^{1/x}}=\left(x^{1/x}\right)^{1/2}={\large x^{1/(2x)}}\), and using some log properties, you have \[\ln L=\lim_{x\to\infty}\frac{\ln x}{2x}=\frac{\infty}{\infty}\] Apply L'Hopital's rule as many times as necessary.

OpenStudy (wiggler):

Thanks!

OpenStudy (anonymous):

yw